# DiffSampling: Enhancing Diversity and Accuracy in Neural Text Generation

### v4

@@ Appendix A Computational Infrastructure and Implementation Details
Appendix B Prompts
@@
Return a summary comprising of 1 sentence. With the sentence in a numbered list format. For example: 1. First sentence assistant
@@
Appendix C Experiments on Temperature Scaling
In addition to investigating performances at a temperature $\tau=1.0$, we also conduct experiments with a lower temperature value ($0.6$) and three higher temperature values ($1.5$, $2.0$, and $10.0$), to verify whether our truncation strategy only preserves appropriate tokens, i.e., whether at different temperatures the quality of generated outputs remains competitive and the diversity increases.
C.1 Math Problem Solving
For the sake of completeness, we also report the full results on a sample of $1000$ entries from the MetaMathQA training set. As apparent from Table 6, the greediness of the approach is directly correlated with the accuracy of solutions. In particular, sampling at a temperature of $0.6$ increases the accuracy of all baselines while undermining their diversity scores, while higher temperatures lower the accuracy and increase (syntactic) diversity; notably, a very high temperature causes semantic diversity to fall. On the contrary, our three methods achieve similar accuracy at any temperature, with small increases in diversity.
C.2 Extreme Summarization
Similar considerations can be traced for XSum, as reported in Table 7. For both RLHF-instructed and pre-trained models, the quality of output produced by the baselines tends to dramatically decrease at higher temperatures (only min-$p$ achieves good results at $\tau=2.0$), with the consequence of an increasing against-greedy diversity due to the choice of random and meaningless tokens. Instead, the quality of the output generated by DiffSampling remains more stable, with small but consistent increases in diversity.
C.3 Divergence Association Task
Figure 4 reports the DAT score and the percentage of output validity of the DPO-instructed and pre-trained models with different temperature values. Top-$p$, locally typical, and $\eta$-sampling rapidly stop outputting valid lists of nouns when the temperature raises, even if the DAT score tends to be higher at $\tau=1.5$; min-$p$ returns a high percentage of valid outputs even at $\tau=2.0$, but increase the DAT score only for the pre-trained model and cannot produce anything valid at $\tau=10.0$. Instead, the performance of our methods remains very similar across different temperatures in terms of both the DAT score and the percentage of valid outputs, except for DiffSampling-lb, whose scores increase at the cost of some validity.
On the other hand, the greedy decoding strategy is less effective for the pre-trained model, which results in higher temperatures yielding better DAT scores across both the baselines and our methods. However, the number of valid outputs decreases faster, and top-$p$, locally typical, and $\eta$-sampling produce very few correct lists at a temperature of $1.5$ (but with a higher DAT score). Again, min-$p$ better manages temperatures around $1.5$ and $2.0$, with higher scores and still at least half of the outputs as valid, but cannot produce any correct output at a temperature of $10.0$.
C.4 WritingPrompts
Appendix D Additional Experiments
D.1 Divergent Association Task over All Nouns
D.2 Temperature Before or After Truncating
D.3 Ablation Study on the Lower Bound
To sum up, when greediness is desirable, a lower value of $p_{lb}$ can lead to high quality and diversity; otherwise, increasing $p_{lb}$ improves diversity, but the cost in terms of validity is not negligible and requires careful consideration.
D.4 Ablation Study on the Dynamic Upper Bound
Table 18 reports the results for the math problem-solving tasks, considering the GSM8K (left side) and MATH (right side) test sets. As expected, the against-greedy diversity scores and cross-input EAD decrease together with $p_{min}$, plateauing at $p_{min}=0.6$ (from that on, results are comparable with DiffSampling-cut); specularly, accuracy is lower at smaller $p_{min}$, but the instructed model reaches a competitive score even at $p_{min}=0.3$. The same holds for XSum as well. As shown in Table 19, diversity decreases when increasing $p_{min}$ and plateaus at 0.5, while quality rapidly increases for the pre-trained model and is almost constant for the instructed model.
To sum up, values above $0.5$ are not different from DiffSampling-cut, while lower $p_{min}$ can help foster diversity with a small loss in accuracy, especially for instructed models.
@@ D.4 Ablation Study on the Dynamic Upper Bound
Method Accuracy Cross-Input Against-Greedy
placeholder placeholder EAD SBERT EAD SBERT
Greedy $95.27_{\pm.17}$ $1.67_{\pm.01}$ $0.71_{\pm.00}$ - -
Temperature = 1.0
Top-$p$ $91.70_{\pm.62}$ $1.71_{\pm.01}$ $0.71_{\pm.00}$ $19_{\pm.01} 0.0$$04_{\pm.00}$
$\eta$-sampling $89.17_{\pm.52}$ $1.73_{\pm.01}$ $0.71_{\pm.00}$ $21_{\pm.01} 0.0$$04_{\pm.00}$
Locally typical $91.70_{\pm.62}$ $1.71_{\pm.01}$ $0.71_{\pm.00}$ $19_{\pm.01} 0.0$$04_{\pm.00}$
Min-$p$ $92.00_{\pm.49}$ $1.70_{\pm.01}$ $0.71_{\pm.00}$ $19_{\pm.01} 0.0$$03_{\pm.00}$
DiffS.-cut $94.93_{\pm.45}$ $1.68_{\pm.01}$ $0.71_{\pm.00}$ $11_{\pm.00} 0.0$$02_{\pm.00}$
DiffS.-lb $92.40_{\pm.50}$ $1.70_{\pm.01}$ $0.71_{\pm.00}$ $19_{\pm.01} 0.0$$03_{\pm.00}$
DiffS.-minp $92.53_{\pm.03}$ $1.70_{\pm.01}$ $0.71_{\pm.00}$ $19_{\pm.01} 0.0$$03_{\pm.00}$
Temperature = 0.6
Top-$p$ $94.03_{\pm.27}$ $1.68_{\pm.01}$ $0.71_{\pm.00}$ $14_{\pm.01} 0.0$$03_{\pm.00}$
$\eta$-sampling $93.90_{\pm.09}$ $1.70_{\pm.01}$ $0.71_{\pm.00}$ $15_{\pm.01} 0.0$$03_{\pm.00}$
Locally typical $94.03_{\pm.27}$ $1.68_{\pm.01}$ $0.71_{\pm.00}$ $14_{\pm.01} 0.0$$03_{\pm.00}$
Min-$p$ $93.93_{\pm.36}$ $1.68_{\pm.01}$ $0.71_{\pm.00}$ $14_{\pm.01} 0.0$$03_{\pm.00}$
DiffS.-cut $94.83_{\pm.18}$ $1.68_{\pm.01}$ $0.71_{\pm.00}$ $11_{\pm.00} 0.0$$02_{\pm.00}$
DiffS.-lb $94.47_{\pm.20}$ $1.68_{\pm.01}$ $0.71_{\pm.00}$ $15_{\pm.01} 0.0$$03_{\pm.00}$
DiffS.-minp $94.17_{\pm.30}$ $1.69_{\pm.01}$ $0.71_{\pm.00}$ $15_{\pm.01} 0.0$$03_{\pm.00}$
Temperature = 1.5
Top-$p$ $87.63_{\pm.90}$ $1.75_{\pm.00}$ $0.71_{\pm.00}$ $24_{\pm.01} 0.0$$05_{\pm.00}$
$\eta$-sampling $80.63_{\pm.23}$ $1.83_{\pm.01}$ $0.71_{\pm.00}$ $28_{\pm.01} 0.0$$05_{\pm.00}$
Locally typical $87.97_{\pm.38}$ $1.77_{\pm.01}$ $0.71_{\pm.00}$ $24_{\pm.01} 0.0$$05_{\pm.00}$
Min-$p$ $87.90_{\pm.57}$ $1.75_{\pm.01}$ $0.71_{\pm.00}$ $24_{\pm.01} 0.0$$05_{\pm.00}$
DiffS.-cut $95.17_{\pm.18}$ $1.67_{\pm.01}$ $0.71_{\pm.00}$ $11_{\pm.00} 0.0$$02_{\pm.00}$
DiffS.-lb $90.73_{\pm.54}$ $1.71_{\pm.02}$ $0.71_{\pm.00}$ $21_{\pm.01} 0.0$$04_{\pm.00}$
DiffS.-minp $91.07_{\pm.10}$ $1.72_{\pm.01}$ $0.71_{\pm.00}$ $21_{\pm.01} 0.0$$04_{\pm.00}$
Temperature = 2.0
Top-$p$ $30.17_{\pm.76}$ $8.30_{\pm.11}$ $0.67_{\pm.00}$ $0.70_{\pm.01}$ $0.44_{\pm.01}$
$\eta$-sampling $42.20_{\pm 1.0}$ $6.26_{\pm.13}$ $0.72_{\pm.00}$ $0.58_{\pm.01}$ $0.29_{\pm.01}$
Locally typical $29.07_{\pm.72}$ $8.68_{\pm.08}$ $0.65_{\pm.00}$ $0.71_{\pm.01}$ $0.45_{\pm.01}$
Min-$p$ $83.93_{\pm.22}$ $1.80_{\pm.02}$ $0.71_{\pm.00}$ $28_{\pm.01} 0.0$$05_{\pm.00}$
DiffS.-cut $94.33_{\pm.35}$ $1.68_{\pm.01}$ $0.71_{\pm.00}$ $12_{\pm.00} 0.0$$02_{\pm.00}$
DiffS.-lb $89.73_{\pm.71}$ $1.72_{\pm.02}$ $0.71_{\pm.00}$ $23_{\pm.01} 0.0$$04_{\pm.00}$
DiffS.-minp $90.13_{\pm.15}$ $1.72_{\pm.01}$ $0.71_{\pm.00}$ $22_{\pm.01} 0.0$$04_{\pm.00}$
Temperature = 10.0
Top-$p$ $0.00_{\pm.00}$ $13.10_{\pm.03}$ $0.12_{\pm.00}$ $1.00_{\pm.00}$ $0.98_{\pm.00}$
$\eta$-sampling $0.00_{\pm.00}$ $13.22_{\pm.03}$ $0.12_{\pm.00}$ $1.00_{\pm.00}$ $0.98_{\pm.00}$
Locally typical $0.00_{\pm.00}$ $13.33_{\pm.02}$ $1_{\pm.00} 1.01±.0$$1_{\pm.00}$ $0.98_{\pm.00}$
Min-$p$ $0.00_{\pm.00}$ $13.21_{\pm.03}$ $0.13_{\pm.00}$ $1.00_{\pm.00}$ $0.98_{\pm.00}$
DiffS.-cut $94.47_{\pm.19}$ $1.67_{\pm.01}$ $0.71_{\pm.00}$ $12_{\pm.00} 0.0$$02_{\pm.00}$
DiffS.-lb $87.17_{\pm.63}$ $1.75_{\pm.01}$ $0.71_{\pm.00}$ $26_{\pm.01} 0.0$$05_{\pm.00}$
DiffS.-minp $87.17_{\pm.81}$ $1.74_{\pm.01}$ $0.71_{\pm.00}$ $25_{\pm.01} 0.0$$05_{\pm.00}$
Table 6: Accuracy and diversity of results for the training set over 3 seeds with different temperature values. The mean and standard error of the final score for each run are reported for accuracy and cross-input diversity, whereas the mean and $95\%$ confidence interval for the full set of answers are reported for against-greedy diversity.

### v5

@@ Appendix B Computational Infrastructure and Implementation Details
Appendix C Evaluation Metrics
In this section, we formally define the quantitative metrics used in our experiments.
  • Cross-Input Diversity: The diversity of outputs across inputs, i.e., the diversity of $\pi(y)$: $D(\bigcup_{i=1}^{N}\pi(y_{i}|x_{i}))$, with $D$ diversity metric (Kirk et al., 2024). In the case of cross-input EAD diversity, this is practically done by computing the EAD score for the entire set of outputs at the same time (i.e., the expected average distinct $N$-grams among all the sequences) for $N\in[1,5]$ and returning their average. We refer to the original paper by Liu et al. (2022) for the exact computation of EAD. On the contrary, in the case of cross-input SBERT diversity, this is practically implemented by first encoding all outputs into latent vectors through a given text encoder, then computing all the possible pair-wise cosine similarity between different vectors, and finally returning $1$ minus the average cosine similarity.
  • Against-Greedy Diversity: The diversity between a given output $y_{i}$ sampled from $\pi(x_{i})$ and the greedy output $y_{i}^{grd}$: $D(y_{i},y_{i}^{grd})$. The against-greedy SBERT diversity requires computing the latent vectors of both outputs and returning 1 minus their cosine similarity. In the case of against-greedy EAD diversity, however, we replace the classic ratio of distinct $N$-gram computation with the ratio of new, distinct $N$-gram from $y_{i}$ with respect to $y_{i}^{grd}$, and then compute EAD for $N\in[1,5]$ starting from it.
  • Per-Input Diversity: The diversity of the output set over a specific input, i.e., the diversity of $\pi(y|x)$: $D(\{y_{i,1}\ldots y_{i,M}\}),y_{i,j}\sim\pi(\cdot|x_{i})$ (Kirk et al., 2024). In the case of per-input EAD diversity, this is implemented by computing the EAD score for the set of outputs from the same input (i.e., the expected average distinct $N$-grams among all the sequences) for $N\in[1,5]$ and returning their average. In the case of the per-input SBERT diversity, this is obtained by first encoding the outputs into latent vectors through a given text encoder, then computing all the possible pair-wise cosine similarity between different vectors, and finally returning $1$ minus the average cosine similarity.
  • Accuracy: The percentage of correctly solved problems, i.e., given the count of solved problems $C_{correct}$ and the count of total problems $C_{total}$, accuracy is defined as $100\cdot\frac{C_{correct}}{C_{total}}$.
  • Rouge-1 (R1): The ratio of $1$-grams present in both the target $y_{i}^{*}$ and the generated output $y_{i}$ (Lin, 2004).
  • Sentence Embedding Cosine Similarity (SIM): The cosine similarity between the latent vector $v_{y_{i}}$ and the latent vector $v_{y_{i}^{*}}$, where $y_{i}^{*}$ is the target output and $y_{i}$ is the generated output. The latent vectors are obtained from a pre-trained text encoder.
  • Coherence (COH): The cosine similarity between the latent vector $v_{s_{i}}$ and the latent vector $v_{y_{i}}$, where $s_{i}$ represents a target text (e.g., a text to summarize) or the input passed to the model, and $y_{i}$ is the output from the model (e.g., the continuation of $s_{i}$ or its summary). The latent vectors are obtained from the SimCSE embedder (Gao et al., 2021).
  • DAT Score: The average cosine distance (i.e., 1 minus cosine similarity) multiplied by $100$ between each pair of word embeddings from a given list of $n$ distinct nouns. The word embeddings are obtained from GloVe (Pennington et al., 2014). $n$ can be either $7$ or $10$. If there are not enough distinct nouns, the score is not computed at all.
  • DAT Valid Outputs: The percentage of experiments in which the count of generated, distinct nouns is greater than or equal to $n$, i.e., the percentage of experiments where it is possible to compute the DAT score.
Appendix D Prompts
@@
Return a summary comprising of 1 sentence. Write the sentence in a numbered list format. For example: 1. First sentence assistant
@@
Appendix E Scaling Model Size
Our main experiments focus only on models of limited size (i.e., 3B, 7B, and 8B models). To demonstrate that our sampling methods can scale up with model size, we also conduct some preliminary tests with a larger model, i.e., with Meta-Llama-3-70B (quantized to 4-bit precision), considering both its pre-trained and its instructed versions. In particular, we repeat all the experiments conducted in Section 5 at a fixed temperature $\tau=1.0$.
Tables 6 and 7 report the results for the math problem-solving case study. As for the smaller model, DiffSampling-cut performs on par or better than greedy in terms of accuracy, while achieving a higher diversity than greedy but lower than the other sampling methods. DiffSampling-lb achieves slightly worse accuracy than top-$p$, but with increases in cross-input EAD. Finally, there are no substantial differences between DiffSampling-minp and min-$p$ apart from cross-input EAD, thus confirming that our relaxation allows for a little more diversity with no cost in terms of quality.
Dataset: RLHF-instructed Pre-trained
Greedy $57.62_{\pm.00}$ $0.75_{\pm.00}$ $0.65_{\pm.00}$ - - $10.16_{\pm.00}$ $0.97_{\pm.00}$ $0.56_{\pm.00}$ - -
Top-$p$ $53.85_{\pm.20}$ $0.88_{\pm.00}$ $0.65_{\pm.00}$ $0.56_{\pm.00}$ $0.11_{\pm.00}$ $7.83_{\pm.14}$ $2.00_{\pm.01}$ $0.59_{\pm.00}$ $0.63_{\pm.00}$ $0.31_{\pm.00}$
$\eta$-sampling $48.37_{\pm.28}$ $0.92_{\pm.00}$ $0.65_{\pm.00}$ $0.60_{\pm.00}$ $0.12_{\pm.00}$ $4.88_{\pm.11}$ $2.22_{\pm.01}$ $0.60_{\pm.00}$ $0.70_{\pm.00}$ $0.33_{\pm.00}$
Locally typical $53.68_{\pm.16}$ $0.88_{\pm.00}$ $0.65_{\pm.00}$ $0.56_{\pm.00}$ $0.11_{\pm.00}$ $7.43_{\pm.21}$ $2.00_{\pm.00}$ $0.59_{\pm.00}$ $0.63_{\pm.00}$ $0.32_{\pm.00}$
Min-$p$ $56.20_{\pm.58}$ $0.85_{\pm.00}$ $0.65_{\pm.00}$ $0.54_{\pm.00}$ $0.10_{\pm.00}$ $9.12_{\pm.93}$ $1.52_{\pm.00}$ $0.56_{\pm.00}$ $0.51_{\pm.01}$ $0.28_{\pm.00}$
DiffS.-cut $59.29_{\pm.63}$ $0.77_{\pm.00}$ $0.65_{\pm.00}$ $39_{\pm.01} 0.0$$08_{\pm.00}$ $12.69_{\pm.20}$ $0.97_{\pm.00}$ $0.57_{\pm.00}$ $0.26_{\pm.01}$ $0.18_{\pm.00}$
DiffS.-lb $52.14_{\pm.60}$ $0.89_{\pm.00}$ $0.65_{\pm.00}$ $0.56_{\pm.00}$ $0.11_{\pm.00}$ $7.25_{\pm.42}$ $2.03_{\pm.01}$ $0.59_{\pm.00}$ $0.64_{\pm.00}$ $0.32_{\pm.00}$
DiffS.-minp $56.00_{\pm.18}$ $0.85_{\pm.00}$ $0.65_{\pm.00}$ $0.54_{\pm.00}$ $0.10_{\pm.00}$ $9.93_{\pm.14}$ $1.56_{\pm.01}$ $0.56_{\pm.00}$ $0.51_{\pm.01}$ $0.28_{\pm.00}$
Table 6: Accuracy and diversity of results for the GSM8K test set over 3 seeds for the instructed (left) and pre-trained (right) Meta-Llama-3-70B model. The mean and standard error of the final score for each run are reported for accuracy and cross-input diversity, whereas the mean and the $95\%$ confidence interval for the full set of answers are reported for against-greedy diversity.
Model: RLHF-instructed Pre-trained
Greedy $22.56_{\pm.00}$ $2.01_{\pm.00}$ $0.80_{\pm.00}$ - - $6.70_{\pm.00}$ $2.18_{\pm.00}$ $0.78_{\pm.00}$ - -
Top-$p$ $20.39_{\pm.15}$ $2.56_{\pm.01}$ $0.79_{\pm.00}$ $0.55_{\pm.00}$ $0.17_{\pm.00}$ $4.04_{\pm.15}$ $5.12_{\pm.02}$ $0.79_{\pm.00}$ $0.57_{\pm.00}$ $0.34_{\pm.00}$
$\eta$-sampling $19.02_{\pm.13}$ $2.80_{\pm.01}$ $0.79_{\pm.00}$ $0.58_{\pm.00}$ $0.18_{\pm.00}$ $3.09_{\pm.09}$ $6.06_{\pm.02}$ $0.79_{\pm.00}$ $0.64_{\pm.00}$ $0.37_{\pm.00}$
Locally typical $20.19_{\pm.20}$ $2.56_{\pm.00}$ $0.79_{\pm.00}$ $0.55_{\pm.00}$ $0.17_{\pm.00}$ $4.12_{\pm.18}$ $5.15_{\pm.02}$ $0.79_{\pm.00}$ $0.57_{\pm.00}$ $0.34_{\pm.00}$
Min-$p$ $20.58_{\pm.04}$ $2.41_{\pm.00}$ $0.79_{\pm.00}$ $0.53_{\pm.00}$ $0.17_{\pm.00}$ $5.46_{\pm.13}$ $3.73_{\pm.01}$ $0.78_{\pm.00}$ $0.45_{\pm.00}$ $0.30_{\pm.00}$
DiffS.-cut $21.27_{\pm.03}$ $2.08_{\pm.00}$ $0.80_{\pm.00}$ $0.40_{\pm.00}$ $0.14_{\pm.00}$ $7.50_{\pm.18}$ $2.29_{\pm.01}$ $0.78_{\pm.00}$ $5_{\pm.00} 0.22±.0$$2_{\pm.00}$
DiffS.-lb $19.79_{\pm.40}$ $2.58_{\pm.01}$ $0.79_{\pm.00}$ $0.55_{\pm.00}$ $0.18_{\pm.00}$ $4.15_{\pm.03}$ $5.22_{\pm.00}$ $0.79_{\pm.00}$ $0.57_{\pm.00}$ $0.34_{\pm.00}$
DiffS.-minp $20.75_{\pm.56}$ $2.45_{\pm.00}$ $0.79_{\pm.00}$ $0.53_{\pm.00}$ $0.17_{\pm.00}$ $5.26_{\pm.21}$ $3.82_{\pm.01}$ $0.78_{\pm.00}$ $0.45_{\pm.00}$ $0.30_{\pm.00}$
Table 7: Accuracy and diversity of results for the MATH test set over 3 seeds for the instructed (left) and pre-trained (right) Meta-Llama-3-70B model. The mean and standard error of the final score for each run are reported for accuracy and cross-input diversity, whereas the mean and the $95\%$ confidence interval for the full set of answers are reported for against-greedy diversity.
The findings from the XSum case study are similar. As reported in Table 8, DiffSampling-cut behaves on par with greedy in terms of quality, but increases scores in terms of per-input diversity. Instead, our two relaxations do not substantially differ from their corresponding baselines top-$p$ and min-$p$. However, it is remarkable that, contrary to what we found for smaller models, greediness does not provide qualitative advantages for the instructed version of Meta-Llama-3-70B.
Model: RLHF-instructed Pre-trained
Greedy $0.29_{\pm.00}$ $0.63_{\pm.00}$ $0.75_{\pm.00}$ $0.20_{\pm.00}$ - - - $0.24_{\pm.00}$ $0.50_{\pm.01}$ $0.70_{\pm.00}$ $0.20_{\pm.00}$ - - -
Top-$p$ $0.29_{\pm.00}$ $0.62_{\pm.00}$ $0.75_{\pm.00}$ $0.50_{\pm.00}$ $0.10_{\pm.00}$ $44_{\pm.01} 0.0$$08_{\pm.00}$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.72_{\pm.01}$ $0.49_{\pm.00}$ $0.71_{\pm.01}$ $0.41_{\pm.01}$
$\eta$-sampling $0.29_{\pm.00}$ $0.62_{\pm.00}$ $0.75_{\pm.00}$ $0.55_{\pm.00}$ $0.11_{\pm.00}$ $49_{\pm.01} 0.0$$09_{\pm.00}$ $0.21_{\pm.00}$ $0.47_{\pm.01}$ $0.61_{\pm.01}$ $0.76_{\pm.01}$ $0.52_{\pm.00}$ $0.75_{\pm.01}$ $0.44_{\pm.01}$
Locally typical $0.29_{\pm.00}$ $0.62_{\pm.00}$ $0.75_{\pm.00}$ $0.50_{\pm.00}$ $0.10_{\pm.00}$ $44_{\pm.01} 0.0$$08_{\pm.00}$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.72_{\pm.01}$ $0.49_{\pm.00}$ $0.71_{\pm.01}$ $0.42_{\pm.01}$
Min-$p$ $0.29_{\pm.00}$ $0.62_{\pm.00}$ $0.75_{\pm.00}$ $0.50_{\pm.00}$ $0.10_{\pm.00}$ $44_{\pm.01} 0.0$$08_{\pm.00}$ $0.23_{\pm.00}$ $0.50_{\pm.01}$ $0.67_{\pm.01}$ $0.65_{\pm.00}$ $0.42_{\pm.00}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$
DiffS.-cut $0.29_{\pm.00}$ $0.63_{\pm.00}$ $0.75_{\pm.00}$ $33_{\pm.00} 0.0$$05_{\pm.00}$ $25_{\pm.01} 0.0$$05_{\pm.00}$ $0.24_{\pm.00}$ $0.50_{\pm.01}$ $0.70_{\pm.00}$ $0.38_{\pm.00}$ $0.18_{\pm.00}$ $0.32_{\pm.01}$ $0.16_{\pm.01}$
DiffS.-lb $0.29_{\pm.00}$ $0.63_{\pm.00}$ $0.75_{\pm.00}$ $0.51_{\pm.00}$ $0.10_{\pm.00}$ $45_{\pm.01} 0.0$$08_{\pm.00}$ $0.22_{\pm.00}$ $0.48_{\pm.01}$ $0.63_{\pm.01}$ $0.73_{\pm.01}$ $0.50_{\pm.00}$ $0.72_{\pm.01}$ $0.42_{\pm.01}$
DiffS.-minp $0.29_{\pm.00}$ $0.63_{\pm.00}$ $0.75_{\pm.00}$ $0.50_{\pm.00}$ $0.10_{\pm.00}$ $44_{\pm.01} 0.0$$08_{\pm.00}$ $0.23_{\pm.00}$ $0.50_{\pm.01}$ $0.66_{\pm.01}$ $0.65_{\pm.00}$ $0.42_{\pm.00}$ $0.63_{\pm.01}$ $0.35_{\pm.01}$
Table 8: Aggregate results over 5 outputs sampled for each of the 1000 prompts from the XSum dataset for the instructed (left) and the pre-trained (right) Meta-Llama-3-70B model. The mean and $95\%$ confidence interval are reported for all the metrics.
The results are similar for the WritingPrompts dataset. As shown in Table 9, the greedy strategy is not optimal for the instructed model; this impacts not only DiffSampling-cut, but also min-$p$ and DiffSampling-minp. However, the other considerations still hold: DiffSampling-cut produces more diverse outputs than the greedy strategy, while the two relaxations perform on par with their corresponding baselines. While greediness appears to guarantee higher quality for the pre-trained model, it should be noted that the per-input EAD becomes extremely low. This is due to extensive repetitions within the same output (but not shared across outputs, which causes the SBERT metric to increase) and may indicate that more greedy strategies perform worse with larger models, which may have learned more unbalanced probability distributions. In such scenarios, looser sampling strategies, such as $\eta$-sampling or our DiffSampling-lb, are likely to be better choices.
Model: RLHF-instructed Pre-trained
Method Quality Per-Input Diversity Quality Per-Input Diversity
placeholder COH EAD SBERT COH EAD SBERT
Greedy $27_{\pm.01} 0.0$$06_{\pm.00} -$ $42_{\pm.01} 0.0$$01_{\pm.00} -$
Top-$p$ $0.29_{\pm.01}$ $0.53_{\pm.00}$ $0.58_{\pm.01}$ $0.26_{\pm.01}$ $0.18_{\pm.01}$ $0.72_{\pm.00}$
$\eta$-sampling $0.29_{\pm.01}$ $0.58_{\pm.00}$ $0.57_{\pm.01}$ $0.26_{\pm.01}$ $0.35_{\pm.01}$ $0.71_{\pm.00}$
Locally typical $0.29_{\pm.01}$ $0.59_{\pm.00}$ $0.46_{\pm.01}$ $0.26_{\pm.01}$ $0.18_{\pm.00}$ $0.72_{\pm.00}$
Min-$p$ $0.27_{\pm.01}$ $0.45_{\pm.01}$ $0.55_{\pm.01}$ $34_{\pm.01} 0.0$$03_{\pm.00}$ $0.60_{\pm.00}$
DiffS.-cut $0.27_{\pm.01}$ $0.23_{\pm.01}$ $0.24_{\pm.01}$ $41_{\pm.01} 0.0$$01_{\pm.00}$ $0.32_{\pm.01}$
DiffS.-lb $0.29_{\pm.01}$ $0.53_{\pm.00}$ $0.57_{\pm.01}$ $0.25_{\pm.01}$ $0.19_{\pm.01}$ $0.72_{\pm.00}$
DiffS.-minp $0.27_{\pm.01}$ $0.44_{\pm.01}$ $0.55_{\pm.01}$ $34_{\pm.01} 0.0$$03_{\pm.00}$ $0.60_{\pm.00}$
Table 9: Aggregate results for the WritingPrompts dataset for the instructed (left) and the pre-trained (right) Meta-Llama-3-70B model. The mean and the $95\%$ confidence interval for the full set of answers are reported for all the metrics.
This consideration finds confirmation from the experiments on the divergent association task. As shown in Fig. 4, DiffSampling-cut always produces the greedy solution for the instructed model, while showing a little more variation for the pre-trained model. Again, both min-$p$ and DiffSampling-minp are influenced by the poor performances of greedy for the pre-trained model, resulting in lower DAT score but higher valid output ratio compared with the other sampling strategies. On the contrary, top-$p$ sampling, locally typical sampling, and DiffSampling-lb seem to better trade off quality and diversity, having a comparably high valid output ratio for the instructed model and especially a substantially higher DAT score for the pre-trained model.
Figure 4: DAT scores for our methods and the baselines over the instructed (left) and pre-trained (right) Meta-Llama-3-70B model. Below, the number of valid outputs produced by each sampling strategy. The dashed line reports the greedy score.
In summary, our preliminary experiments on a larger model confirm the positive relation between our three methods and the corresponding baselines, while also demonstrating that models with different sizes may necessitate different degrees of greediness and thus suggesting a looser sampling strategy (e.g., DiffSampling-lb) when dealing with sufficiently large language models.
Appendix F Experiments on Temperature Scaling
In addition to investigating performance at a temperature of $\tau=1.0$, we conduct experiments at a lower temperature ($0.6$) and three higher temperatures ($1.5$, $2.0$, and $10.0$) to verify whether our truncation strategy preserves only appropriate tokens, i.e., whether the quality of generated outputs remains competitive and diversity improves across different temperatures. Overall, we found that DiffSampling-cut maintains the same level of quality even at very high temperature, and, thus, can be safely used even with $\tau\gg 1.0$. On the other hand, DiffSampling-lb tends to produce lower-quality outputs at higher temperatures. This effect is particularly pronounced for non-instructed models, where performance for $\tau>1.0$ decreases rapidly. However, fine-tuned models maintain competitive quality scores at higher temperatures while increasing diversity; the choice of $\tau$ is a matter of trading off quality and diversity. When accuracy is not a sine qua non requirement, we suggest adopting a $\tau\geq 2.0$. Finally, DiffSampling-minp exhibits the same general trend as DiffSampling-lb for pre-trained models, although the effect is less pronounced. For tuned models, however, it shows virtually no performance degradation at higher temperatures, enabling the use of larger values of $\tau$ in most scenarios.
F.1 Math Problem Solving
@@
F.2 Extreme Summarization